Find the PDF of W = X + Y when X and Y have the joint PDF fx,y(x,y)=ĺ0 x,y(x,y) = otherwise otherwise
Answer #1

# Exy Co, y) = Sto OS X Lys & otherwise Fwcwl be the Let x+y=w., & cod f. of w then ty-w P (WEW) For (W) = - - P(X+YEW) P(x <4 y = 1-s t de dy w-y 4 11 -wtys dy O M = 1- to (ay-w, dy =t- [ % (yz wy)] =1((16-46) - 5 ( .492)] = - ( 16-46 +12) = 1- to (Now. p.dif of w is fw (W) = a FW (W) ocwe4 310 16-20 32 4swed wy P ocwc4 f (W) = 4<wka otherwise